Ticket Number: HEA-NOC/20030319-02 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <=> JANET
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20030320 10:30 UTC
Ticket Closed: 20030407 09:11:47 UTC Problem End: 20030320 17:30 UTC
Retermination of Dublin - Belfast line to JANET PoP in Citywest and bringing up of BGP peering on secondary local GE circuit.
No traffic should be affected as all traffic should now be flowing over the primary local GE and the IE-UK circuit. Equally no traffic should flow over the secondary GE circuit due to BGP metrics.
20030320 13:15 BN
Belfast - Dublin circuit has been reterminated and JANET will configure this link. Bringing up BGP peering on secondary circuit is due to go ahead c. 15:00
20030320 15:50 BN
Secondary GE link has now established a BGP session with JANET. Their filters are still in place so no routes are being sent. These filters will either be brought down later this afternoon or early next week. No traffic should flow over this link.
20030327 13:30 BN
Filters on second local circuit will be removed by JANET Noc on Monday 31st at 10:00 UTC +1. No traffic should flow over his circuit due to MEDs.
20030407 10:15 BN
This circuit is now up and stable. MEDs are working properly, no traffic is flowing over the backup circuit. Ticket being closed.
Time to Fix: